A32 year old mother realized that her age is twice the sum of her two children's age. Her son jose is two years older than.her daughter sophia.
Answer by macston(5194)

.
S=Sophia's age; J=Jose's age=S+2yrs
2(S+J)=32 years Substitute for J
2(S+S+2yrs)=32 years
2(2S+2yrs)=32 years
4S+4yrs=32 years Subtract 4 years from each side.
4S=28 years Divide each side by 4
S=7 years ANSWER 1: Sophia is 7 years old.
J=S+2yrs=7yrs+2yrs=9yrs ANSWER 2: Jose is 9 years old.
.
CHECK:
2(S+J)=32 years
2(7yrs+9yrs)=32 years
2(16yrs)=32 years
32 years=32 years